Suppose that the residents of Windhoek spend all of their income on cauliflower, broccoli and carrots. In 2019 they buy 100 heads of cauliflower for N$2000, 50 bunches of broccoli for N$750 and 500 carrots for N$500. In 2020 they buy 75 heads of Cauliflower for N$2250, 80 bunches of broccoli for N$1200 and 500 carrots for N$1000. Suppose that the base year is 2019.


(a)What is the CPI in both years?(8)

(b) What is the inflation rate?( 4)

(c) Find nominal and real GDP for 2019 and 2020(6)

(d) Find nominal growth rateand real growth rate between 2019 and 2020.(4)

(e ) Which do you think has a greater effect on the CPI: a 10 percent increase in the price of chicken or a 10 percent increase in the price of cars? Why?(3)


1
Expert's answer
2022-03-07T10:43:41-0500

a)

CPI is known to measure the cost of the market basket in the current year and base year and is the ratio of the cost of the market basket in the current year to the base year times 100.


"CPI_{2019}= \frac { 2,000+750+500}{2,000+750+500}\times 100"


"= 1\times 100"


"=100"


"CPI_{2020}= \frac {2,250+1,200+1,000}{{20\times 75}+{15\times 80}{1\times 500}}\times 100"


"=\frac {4,450}{1,500+1,200+500}\times 100"


"=\frac {4,450}{3,200}\times 100"


"=139.06"


b)

inflation rate


"Inflation \space Rate=\frac{CPI_{2020}-CPI_{2019}}{CPI_{2019}}\times 100"


"= \frac {139.06-100}{100}\times 100"


"= 39.06\%"


c)


"Nominal \\space GDP_{2019\u200b}=Real \\space GDP_{2019\u200b}=\\$2,000+\\$750+\\$500=\\$3,250"


"Nominal\\space GDP_{2020\u200b}=\\$2,250+\\$1,200+\\$1,000"


"=\$4,450"


"Real\\space GDP_{2020\u200b}=\\$20\\times 75+\\$15\\times 80+\\$1\\times 500"


"=\$1,500+\$1,200+\$500"


"=\$3,200"


d)


"Growth \space rate=\frac {GDP\space in\space 2020-GDP \space in \space 2019}{GDP \space in \space 2019}"


"=\frac {4,450-3,250}{3,250}\times 100"


"=0.3692\u00d7100"


"= 36.92\%"


Hence the growth rate is 36.92%


e)

since it is evident that chicken has a larger component of the average consumer's market basket, a rise of 10% in the price has a greater impact on the consumer price index than a 10% increase in the price of a vehicle and car is not part of the consumer basket, Therefore, the increase in car price will not show much effect on the CPI.